OpenStudy (anonymous):

Need help now my dad and I are racing Events A and B are mutually exclusive. Also, P(A) = 0.15 and P(B) = 0.2. What is P(A or B)? A. 0.35 B. 0.03 C. 0.20 D. 0.05

OpenStudy (anonymous):

so it A?

OpenStudy (anonymous):

if A and B are mutually exclusive, we're being told that there is no possibility that both occur together, i.e. \(P(A\cap B)=P(A|B)=P(B|A)=0\). now, what does this tell us about \(P(A\cup B)\)?

OpenStudy (anonymous):

recall \(P(A\cup B)=P(A)+P(B)\color{red}{-P(A\cap B)}\)... and we're told that \(P(A\cap B)=0\), so \(P(A\cup B)=P(A)+P(B)\color{red}{-0}=P(A)+P(B)\)